# MOGO: Residual Quantized Hierarchical Causal Transformer for High-Quality and Real-Time 3D Human Motion Generation

### v1

@@ Appendix B Ablation Study
| Epochs | w/ 0% CMP | w/ 50% CMP | w/ 100% CMP |
| 50 | 0.671 | 0.601 | 0.501 |
| 80 | 0.437 | 0.379 | 0.284 |
| 100 | 0.347 | 0.263 | 0.205 |
Input Condition. We compared the performance of input conditions by either adding the prompt token and layer token as a sequence prefix or not. As shown in Table 7, our experiments reveal that when without PnQ (adding prompt token to quantization layer token) is applied, the correlation between generated motion sequences and text (measured by R-Precision and MM Distance) decreases significantly. This may result from an additional layer token between action sequence tokens and prompt tokens during training, leading to some attention loss, as shown in Figure 10.
| Codebook size | w/ PnQ | ||
| FID | Top1 | MM-Dist | |
| 2048 512 | 0.0520.005 | 0.4980.002 | 2.9970.008 |
| 4096 256 | 0.0470.005 | 0.5010.002 | 2.9010.008 |
| 8192 128 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
| Codebook size | w/o PnQ | ||
| FID | Top1 | MM-Dist | |
| 2048 512 | 0.0960.005 | 0.4520.002 | 3.2930.009 |
| 4096 256 | 0.0910.005 | 0.4590.002 | 3.1410.007 |
| 8192 128 | 0.0850.002 | 0.4820.003 | 3.1270.008 |
Appendix C Algorithm
The core training code structure of our MoSA-VQ model is illustrated below. The variable abbreviations used in the code are defined as follows:
The core training code structure of our RQHC-Tranformer is illustrated below. The variable abbreviations used in the code are defined as follows:

### v2

@@ C-E Effect of Codebook Size
We conduct an ablation study to investigate how different codebook sizes affect both reconstruction and generation quality. Specifically, we vary the number of entries and dimensions in the codebook, and evaluate performance on the HumanML3D test set in terms of FID, Top-1 Accuracy, and MM-Distance. Results are summarized in Table VII.
As shown in the table, larger codebooks generally lead to better performance, particularly in the generation setting. When increasing the codebook size from to , the generation FID steadily decreases from 0.203 to 0.038, indicating a substantial improvement in motion realism. Similarly, Top-1 Accuracy improves from 0.469 to 0.515, and MM-Distance also decreases, suggesting enhanced semantic alignment and diversity.
In the reconstruction setting, we observe a similar trend: the FID drops from 0.022 to 0.013 as the codebook size increases. Notably, while some intermediate configurations (e.g., ) achieve marginally better reconstruction FID (0.015), they underperform in generation quality, highlighting that optimal reconstruction does not always translate to optimal generation.
The best overall performance is achieved with a codebook size of , which provides a favorable balance between code granularity and diversity. This configuration yields the lowest FID in both reconstruction and generation settings, while maintaining competitive Top-1 Accuracy and MM-Distance. These results demonstrate that increasing the expressiveness of the codebook—by enlarging its capacity—enables the quantizer to better capture fine-grained motion semantics and contributes to improved generative performance downstream.
C-F Effect of Cross-Level Decorrelation Loss
| decorrelation loss | FID | R TOP1 | MM-Dist |
| w/o | 0.044±0.001 | 0.503 ±0.001 | 2.996 ±0.006 |
| w/ | 0.038±0.001 | 0.508 ±0.001 | 2.992 ±0.008 |
To validate the effectiveness of our proposed cross-level decorrelation loss, we conduct an ablation experiment on the HumanML3D test set, as shown in Table VIII. We compare two settings: one using the full training objective (including ), and one without it.
When the decorrelation loss is removed, we observe a slight degradation in generation quality, with FID increasing from 0.038 to 0.044. Moreover, R-Top1 accuracy drops from 0.508 to 0.503, and MM-Distance increases from 2.992 to 2.996, indicating reduced semantic alignment and less precise generation.
This performance gap, although modest in absolute value, is consistent across all metrics and can be attributed to the inter-level redundancy present when is absent. As defined in Eq. (3), our decorrelation loss minimizes the batch-wise covariance between the quantized vector and the residual at the next level. This enforces statistical orthogonality across quantization layers, effectively disentangling the hierarchical representations.
From an information-theoretic standpoint, this encourages each level to capture novel, non-overlapping features, which reduces redundancy and improves latent code efficiency. In practice, this regularization leads to more compact and expressive codes—especially important under constrained codebook budgets or when deeper quantization hierarchies are used.
In summary, the proposed cross-level decorrelation loss plays a crucial role in enhancing both the semantic clarity and compression efficiency of our quantization framework, and leads to measurable improvements in motion generation quality.